Preparation for Machine CheckOut Status - PowerPoint PPT Presentation

Loading...

PPT – Preparation for Machine CheckOut Status PowerPoint presentation | free to download - id: a2cb6-NDk3O



Loading


The Adobe Flash plugin is needed to view this content

Get the plugin now

View by Category
About This Presentation
Title:

Preparation for Machine CheckOut Status

Description:

Definition of possible dates for preparation tests. Interfaces with other systems ... page connected to LHCCWG and Dry runs pages Reyes once she is back from holidays ... – PowerPoint PPT presentation

Number of Views:33
Avg rating:3.0/5.0
Slides: 19
Provided by: gianluig
Learn more at: http://lhccwg.web.cern.ch
Category:

less

Write a Comment
User Comments (0)
Transcript and Presenter's Notes

Title: Preparation for Machine CheckOut Status


1
Preparation for Machine Check-OutStatus
  • Gianluigi for the Machine Check-Out Team
  • (Mike, Reyes, Rossano)
  • .and many thanks to Georges-Henry, Jörg, Markus
    and many others

2
Approach.so far
  • Extended LTC presentations used as starting point
    for the discussion and to address the (many) open
    questions or unclear points
  • Identification of the Machine Check-Out contacts
    (possibly people on the field) for each System
    (largely based on Rogers table at OP WS 11/07)
  • First round of discussions to collect
    information
  • Plans for the HW/SW availability
  • Status of the commissioning
  • Requirements for commissioning time in
    preparation for beam
  • Status of the tools in CCC (Alarms, logging,
    Application SW) and feedback on them
  • Definition of possible dates for preparation
    tests
  • Interfaces with other systems
  • What needs to be done during a standard cold
    check-out
  • For the moment the contact is very useful and
    informative but the information flow is still
    slow (people busy on the optics of finalizing
    their HW)

3
Systems Table
üContacted - ûNot contacted, yet MsMrsMr-X
eqpt. contact
4
Systems Table
5
Access DSO tests
  • Non-conformities found in week 2 being fixed.
  • LHC DSO tests proposed dates (after discussion
    with TS/AB/HWC to be further discussed with the
    experiments, first somehow expected- reaction is
    negative)
  • Preparation by TS/ASE starting 5/6/2008 1200.
  • Global DSO test 1 6-9/6/2008 (included) in
    addition to the TI2 DSO test. Patrol on 6/6/2008.
  • Global DSO test 2 20-23/6/2008 (included).
    Patrol on 20/6/2008.
  • Need to go through committees like ICC, etc. to
    try and get a rubber stamp (?)
  • They add (and partly overlap) to the TI2 and TI8
    DSO tests and beam commissioning (organized by
    Rossano)
  • TI8 DSO test 16/5/2008
  • TI8 beam test 24-25/5/2008
  • TI2 DSO test during global DSO test 1
  • TI2 beam test 14-15/6/2008

6
Beam Instrumentation
  • See JJs presentation at LHCCWG 11/3/08
  • Activities supposed to start 10/4 ? not yet..
  • Need to start ASAP (JJ promised this week!!) ? By
    middle of May Cool-down of sector 4-5 ? access
    will then be difficult for the instrumentation in
    LSS4
  • BTV system in a pretty good shape (used for Dry
    runs) ? although not all BTVs under control, yet
  • Matching monitor after discussions at the e-LTC
    agreed with BI to have turn-by-turn acquisition
    on BSRT (in parallel to standard slow
    acquisition) from day 1 with possibility of
    installing a fast camera on one of the OTR
    screens in LSS4 (2-4 hours access) in case of
    problem (the infrastructure for that is being
    prepared).
  • BLMs are certainly a major/complex system. Coming
    up with logging in some points. EMC tests done
    with kickers (in 2, 8 and 6, need to do in point
    4). Not yet with magnets (suspected interference
    with TL magnet pulsing tbc). Need reliability
    run and verification of concentrator and logging
    ? Discussion with Bernd

7
BIC
  • Issue of the installation of the optical fibres
    seems solved should be completed by the end of
    April
  • Preliminary planning for the IST
  • trying to discuss the order with Bruno in order
    to get a better match with availability of the
    sectors from HWC
  • Overall cold check-out requirements
  • Some useful information in MPS Commissioning
    Procedures
  • to be detailed and quantified ? Jörg and Bruno
    (with Jan and Alick).
  • First version of the BIC supervision programme
    available (experience from SPS). Alarms will
    appear there.

8
Collimators
  • All HW installed (except TCTVB in pt. 2 and 8 due
    to contamination of the tank)
  • IST ongoing (expected to be finished by end of
    May). Dry runs started ? Mike, Stefano
  • Overall cold check-out requirements
  • The tests required to establish Machine
    Protection Functionality (i.e. tests during cold
    check out with all the other systems) only
    sketched ? Ralph promised that for this week
    (also requested by MPSC)
  • Logging being put in place (Stefano)
  • Alarms need to be defined (Georges-Henry, Markus,
    Stefano)
  • First version of the fixed displays available
  • Protection with RBAC and MCS to be tested
  • Interface with BLMs (BLM concentrator, fast BLM
    acquisition to be tested)
  • Separation between collimators and injection/dump
    protection devices for the controls not totally
    clear

9
Cryogenics
  • From the HWC experience main issue is the
    Cryogenics support (resources for operation)
  • Tools to get the status of the cryogenics are/are
    being made available
  • Cryo-start/maintain status for each sector
  • GTPM view with dependencies from other systems
  • Normalized heat loads (initially normalized cryo
    -valve opening and temperatures, later normalized
    virtual heat-loads) per cell ? useful to identify
    possible localized losses, vacuum bumps
  • Insulation vacuum warnings might be useful ?
    being discussed
  • No time required for Machine Check-Out. Although
    having the above displays soon might be useful to
    follow the cryo behaviour as soon as Sectors
    become available

10
Power Converters
  • No major issues most of the tests done during
    HWC, nevertheless
  • EMC tests not completed (in particular with
    kickers)
  • Exercising with nominal currents and trimming
    (e.g. behaviour close to zero current)
  • Need to be done before taking beam ? possibly as
    we take the sectors coming-out of the HW
    commissioning
  • Alarms are available (need to verify reduction
    criteria and details)
  • Logging implemented (to be verified
    systematically)

11
QPS/EE
  • From the HWC experience main issue is the expert
    support (resources for operation)
  • No additional tests are required during cold
    check-out
  • Tests will be required after every SD ? K.
    Dahlerup-Petersen, R. Denz to provide the
    procedures for future reference
  • Alarms (available for some of the sectors, not
    yet for 5-6)

12
RF - Transverse Feedback - Beam Observation
  • All HW installed in the tunnel
  • In general the power part is in a good shape
  • Low-level part and even more control SW are
    lagging behind
  • Fast signals (pre-pulses, frev, 40 MHz to expt.,
    etc.) not yet distributed
  • The same for observation tools (FESA classes and
    application SW, OASIS signals not available)
  • Expecting a summary of the situation and schedule
    (by the end of this month)
  • Alarms and logging available. But only for RF
    cavities for the time being
  • No specific requests of time for Machine
    Check-Out, so far

13
Timing
  • HW installed and running but
  • Safe Machine Parameter Controller (carrying
    information about energy, intensity) not yet in
    its final version ? Simulated values
  • Beam Permit flags not yet wired to timing
  • Schedule to be discussed with B. Puccio
  • Monitoring tools for the timing is very important
    as timing is playing a critical role (more than
    in the SPS or LEP) in the safe operation of the
    machine
  • Alarms when timing events are not sent or delayed
  • Tools to dump expected vs. distributed events are
    needed in case of error ? Julian could have a
    tool for that
  • All the mechanics of the injection must be tested
    all through the chain both for the pilot filling
    scheme for commissioning and for the nominal
    filling scheme ? Dry run on 8/5 being organized.

14
Vacuum
  • Vacuum installation almost complete. Expect to
    close and pump the experimental areas by the end
    of May. Still true? Waiting for plan from AT/VAC
  • Vacuum valves and their control tested but for
    the moment not in remote control (disconnected
    and blocked IN) ? need the beam info from BIC
  • Alarms are available for beam/insulation vacuum ?
    reduction to be applied to avoid too many alarms
    when problems ? Markus, Georges-Henry.
    Information about insulation vacuum could be
    useful to anticipate possible problems for the
    cryogenics
  • Logging
  • defined for the vacuum pressures (both for
    insulation and beam vacuum but for the latter no
    data yet) with an acquisition rate of lt1/min.
    Sufficient for the beam pressures?
  • for position of the valves on change.
  • Application is OK some minor modifications
    requested ? presentation at the next LHCCWG (I.
    Laugier)

15
General Issues
  • Verification of UPS connection of critical
    equipment is (was) part of the HWC not done for
    sector 4-5 and 5-6 (so far) ? impact on
    cryogenics and organization
  • Status of the compensators and behaviour at full
    load (e.g. Pt. 2 and Pt. 4)
  • Post-mortem not really addressed yet
  • Logging alarms ? entries are coming ? needs
    some order and verification
  • Trying to cope with the little solid
    information on the overall HWC and Installation
    plans for the time beingeven more difficult now
    that HWC is not appearing in the official
    planning
  • Issue of the DSO tests dates ?

16
Next step take ownership of Sector 5-6
  • Started discussions with HWC (second week of
    May). Useful for the definition of the procedures
    of the cold check-out
  • Generate and load nominal functions 5 TeV ramp
    all the magnets
  • Training in the late afternoon (only dipoles?)
  • Trim (chromaticity, tune, coupling, beta beat,
    orbit correction) and verify the effect of the
    trim.
  • Connect BIC (beam info needed for the vacuum
    valves), t.b.c.
  • Vacuum controls in remote (valves, vacuum pumps)
  • Add distributed instrumentation (BLMs already
    available in L6-, BPMs, BTVs, RAMSES, RADMON)
  • Add Beam dump elements
  • Add collimators (few) and protection elements
    (available?)
  • SIS verification
  • Verify fixed displays, alarms (by sector),
    logging

17
Hand-over of a Sector
  • We need to know/have from HWC
  • List of tests done
  • List of maximum currents
  • List of non-conformities. There are a lot of them
    with decision pending. Shall we go through them
    here before taking over a sector? Can we do
    differently?
  • Access in the machine in the handed-over sectors
    should go through the EIC on shift
  • .anything else?
  • Needs to run handed-over sector
  • Availability of a (at least) a console for
    Machine Check-Out
  • Some Wall displays
  • Suggestion TI corner close to the Cryo people

18
To be put in place
  • Names of people attached to each of the system
    tests for 5-6
  • I guess that Mike, Jörg, Reyes, Rossano and
    myself will be around most of the time, but
  • Tests of the various systems
  • Applications
  • Verification of the logging
  • Alarms
  • Should be done with the help of the system
    commissioners (e.g. vacuum ? Frank,) on agreed
    periods, as well as the follow-up of specific
    issues
  • Information collected so far and follow-up
    first embryo of the check-out procedures will be
    put in a web-repository/page connected to LHCCWG
    and Dry runs pages ? Reyes once she is back from
    holidays
  • Starting on the middle/end of May
  • Should start more regular meetings to review
    progress in the preparation of the HW with eqpt.
    contacts. (merging with ABCIS?)
  • Regular reports to LHCCWG
About PowerShow.com